OpenStudy (anonymous):

HELP! The probability that Roger wins a tennis tournament is 0.75 (event A), and the probability that Stephan wins a golf tournament is 0.64 (event B). The probability that both Roger and Stephan win their respective tournaments is 0.48. Which statement is true? Events A and B are independent because P(A and B) = P(A) P(B). Events A and B are independent because P(A and B) = P(A) + P(B). Events A and B are dependent because P(A and B) = P(A) P(B). Events A and B are dependent because P(A and B) = P(A) + P(B). Events A and B are dependent because P(A and B) P(A) P(B).

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

P(A) = 0.75 P(B) = 0.64 P(A & B) = 0.48 P(A) * P(B) = 0.75*0.64 = 0.48 = P(A & B) Which means A and B events are independent. ANS: Events A and B are independent because P(A and B) = P(A) P(B).

OpenStudy (madhu.mukherjee.946):

Events A and B are independent because P(A and B) = P(A) P(B).

OpenStudy (anonymous):

thank you both so much!

OpenStudy (madhu.mukherjee.946):

your welcome